when a number is decreased by 7 and this result is divided by the sum of the number and 11, the answer is -3 over 4. find the number.
Answer by stanbon(75887)

when a number is decreased by 7 and this result is divided by the sum of the number and 11, the answer is -3 over 4. find the number.
---------------------
Let the number be "x".
Equation:
(x-7)/(x+11) = (-3/4)
Cross-multiply:
4(x-7) = -3(x+11)
4x-28 = -3x-33
7x = -5
x = -5/7
===============
Cheers,
Stan H.